# Dual-VQE: A quantum algorithm to lower bound the ground-state energy

### v2

@@ III.2.1 Pretraining using matrix product states
The dual-VQE problem (without the final parameterization of the state $\omega$) is given by
@@
We begin with a randomly initialized MPS with $n_{S}+n_{R}$ nodes in the canonical form, where $n_{R}$ is the number of qubits in the reference system $R$, and $n_{S}$ is the number of qubits in the system $S$ of interest. Again, we pick $n_{S}$ to be at least as large as $n_{R}$ to ensure a full-rank mixed state on the system qubits. The two terms that depend on $\omega$ are $\operatorname{Tr}[H\omega]$ and $\operatorname{Tr}[\omega^{2}]$. These terms are estimated using the tensor networks in Figure 1b) and Figure 1c), respectively.
@@
III.2.2 Convert final matrix product state to parameterized quantum circuit
After pretraining, we convert the resulting MPS into a parameterized quantum circuit composed of one- and two-qubit gates. If the MPS has a maximum bond dimension of two ($\chi=2$), it is exactly decomposable into a single layer of two-qubit unitaries [45]. However, for an arbitrary MPS with $\chi>2$, an exact decomposition into two-qubit gates is not generally feasible. This is because each MPS bond with dimension $\chi>2$ corresponds to a unitary acting on $\lceil\log_{2}(\chi)\rceil+1$ qubits, and it is not always possible to decompose any multi-qubit unitary exactly into a sequence of two-qubit gates.
To translate the final MPS into a parameterized quantum circuit, we make use of the algorithms put forth in [46]. The building blocks of the different algorithms are called the Analytical Decomposition (AD) algorithm and Optimizing Decomposition (OD). We note that in this work and in prior works, the AD algorithm is also called the disentangling algorithm.
@@ III.2.2 Convert final matrix product state to parameterized quantum circuit
While the translated circuit is now composed of only two-qubit unitaries, it still needs to be cast into a form realizable on a quantum device. To that end, we utilize the KAK decomposition to decompose each two-qubit unitary into a product of parameterized rotation gates [49]. Each two-qubit unitary contributes a total of 15 parameters.
III.2.3 Train parameterized quantum circuit until convergence
Lastly, we use the translated MPS to warm-start the quantum optimization. Before we begin the quantum training, we append additional layers of the parameterized ansatz that are trivially initialized (identity operations). Then, all the parameters, including those from the pretrained portion of the circuit, are trained.
@@ IV.1 Purification and convex combination without pretraining
Prior to training, we initialized the parameters $\eta$, $\mu$, and $\theta$. We set the parameters $\eta$ and $\nu$ in (13) to be equal to zero and one, respectively. We selected the components of the initial parameter vector $\theta$ for the state $\omega(\theta)$ uniformly at random from $0,2\pi]$. The optimal state $\omega$ is almost always a full-rank mixed state. Thus, we used either a three-layer purification ansatz or a convex-combination ansatz with a two-layer quantum circuit Born machine and a two-layer unitary. The particular structure of the purification ansatz and the quantum circuit Born machine we use can be found in a figure in our companion paper [8].
@@
The variation displayed in the graphs comes from three sources of randomness: the randomized initializations of all parameterized quantum circuits, randomness in the SPSA algorithm for gradient estimation, and shot noise over $10^{12}$ shots for every measurement of a quantum circuit. In our simulations, we were able to achieve an error of order $10^{-2}$ after $20,000$ iterations of training. These results give evidence that using VQE and dual-VQE allows us to bound the true value from both above and below, respectively, as desired.
@@ IV.2 MPS pretraining with purification ansatz
Figure 3: Objective function and penalty value for the Ising model problem using dual-VQE and MPS pretraining. After 2,000 iterations of pretraining followed by 25,000 iterations of quantum training, the true solution is reached when $\chi_{\max}=8$, with a relative error of 0.5%.
@@
The results of the different experiments where we used MPS pretraining are given in Figure 3. The plots shown are for $n=3$ qubits. To demonstrate the effectiveness of pretraining, we used three different max bond dimension values $\chi_{\operatorname{max}}=\{2,4,8\}$ and compare it with the no-pretraining case. In each of the above cases, we conducted five runs and report the mean and standard deviation.
@@
Following the pretraining, we translated the MPS into a quantum circuit. We used the AD decomposition to decompose the MPS into three layers of two-qubit unitaries, labelled as Disentangler in Figure 4. We plot the fidelity of translation, using the AD and OD algorithms (defined in Appendix B), in Figure 4.
The circuit we obtained then serves as the starting point for one of the OD algorithm runs. For this run, labelled as Warm Start, we used 2000 iterations and a constant learning rate of $\beta=0.2$ (see (31)). We also ran the OD algorithm on a randomly initialized three layer circuit and trained for 2000 iterations, labelled as Random Start. Finally, we picked the best translated circuit. For $\chi_{\max}=8$, the translation fidelity was around $99.5\%$.
Lastly, we added a single layer of parameterized gates to the translated circuits, as described in Section III.2.3, and all the parameters are trained using a combination of a finite difference method and the SPSA algorithm. For the parameters $\eta$ and $\nu$, we used a finite difference method with a perturbation of $10^{-6}$. For the circuit parameters, we used the SPSA algorithm since it provides an unbiased estimate of the gradient with a runtime complexity that is independent of the number of parameters. For $\eta$ and $\mu$, we initialized the learning rate, denoted by $\beta_{\eta,\mu}$, to be $10^{-3}$ and for the circuit parameters, we initialized the learning rate, denoted by $\beta_{\omega}$, to be $3\times 10^{-3}$. Every 200 iterations, we modified the learning rates and the penalty parameter $c$. For the learning rates, we calculated the slope of the last 200 iterations. If the slope was negative, we set
@@
$\beta_{\eta,\mu}$ $=\max(0.9\beta_{\eta,\mu},3\beta_{\min}).$ (35)
@@
$\beta_{\eta,\mu}$ $=\min(1.02\beta_{\eta,\mu},3\beta_{max}).$ (37)
@@ Appendix A Matrix product states
Matrix product states have a linear architecture, with each tensor being connected to their nearest neighbours. For example, the tensor $A^{\psi}$ can be factorized into a linear chain of smaller tensors, as seen in Figure 6.
@@
where we use the Einstein repeated index summation convention. The indices labelled by $b_{k}$ are called bonds, and their dimension is called the bond dimension of the MPS. Note that, in general, the bonds can be of different dimension, in which case, the largest bond is quoted as the bond dimension of the MPS and is denoted by $\chi$. In addition to the bond dimension, the size of the output indices is called the physical dimension. In this work, since we are working with qubits, we fix the physical dimension to be $2$. Thus, the number of parameters of the entire MPS is given by $2n\chi^{2}$, as opposed to $2^{n}$ for the full state vector.
@@
The advantages of using tensor networks do not end here. Matrix product operators generalize unitary operations and can be implemented using tensor contractions in a runtime polynomial in the bond dimension. For states with entanglement greater than the max bond dimension can exactly reproduce, the singular value decomposition may be used to find a low-bond dimension MPS approximation [56].
Just as mixed states in a Hilbert space can be represented by a pure state in a larger Hilbert space, so too can larger rank MPSs be used to represent mixed states on a smaller rank. For example, consider a mixed state $\omega_{S}$ that has a purification $|\psi\rangle_{RS}$. If $|\psi\rangle_{RS}$ has an MPS representation, then $\omega_{S}$ can be realized as an MPO using this MPS, as depicted in Figure 1a).

### v3

@@ III.2.1 Pretraining using matrix product states
The dual-VQE problem (without the final parametrization of the state $\omega$) is given by
@@
We begin with a randomly initialized MPS with $n_{S}+n_{R}$ nodes in the canonical form, where $n_{R}$ is the number of qubits in the reference system $R$, and $n_{S}$ is the number of qubits in the system $S$ of interest. Again, we pick $n_{R}$ to be at least as large as $n_{S}$ to ensure a full-rank mixed state on the system qubits. The two terms that depend on $\omega$ are $\operatorname{Tr}[H\omega]$ and $\operatorname{Tr}[\omega^{2}]$. These terms are estimated using the tensor networks in Figure 1b) and Figure 1c), respectively.
@@
III.2.2 Convert final matrix product state to parametrized quantum circuit
After pretraining, we convert the resulting MPS into a parametrized quantum circuit composed of one- and two-qubit gates. If the MPS has a maximum bond dimension of two ($\chi=2$), it is exactly decomposable into a single layer of two-qubit unitaries [42]. However, for an arbitrary MPS with $\chi>2$, an exact decomposition into two-qubit gates is not generally feasible. This is because each MPS bond with dimension $\chi>2$ corresponds to a unitary acting on $\lceil\log_{2}(\chi)\rceil+1$ qubits, and it is not always possible to decompose any multi-qubit unitary exactly into a sequence of two-qubit gates.
To translate the final MPS into a parametrized quantum circuit, we make use of the algorithms put forth in [43]. The building blocks of the different algorithms are called the Analytical Decomposition (AD) algorithm and Optimizing Decomposition (OD). We note that in this work and in prior works, the AD algorithm is also called the disentangling algorithm.
@@ III.2.2 Convert final matrix product state to parametrized quantum circuit
While the translated circuit is now composed of only two-qubit unitaries, it still needs to be cast into a form realizable on a quantum device. To that end, we utilize the KAK decomposition to decompose each two-qubit unitary into a product of parametrized rotation gates [51]. Each two-qubit unitary contributes a total of 15 parameters.
III.2.3 Train parametrized quantum circuit until convergence
Lastly, we use the translated MPS to warm-start the quantum optimization. Before we begin the quantum training, we append additional layers of the parametrized ansatz that are trivially initialized (identity operations). Then, all the parameters, including those from the pretrained portion of the circuit, are trained.
III.2.4 Discussion of scalability
In Section IV, we present simulated results for a small example problem of interest and find that the MPS pretraining procedure outlined above speeds up the quantum training significantly. Here, we appeal to prior numerical results [45] to argue that this performance advantage is expected to persist even as we increase the number of qubits beyond the reach of classical brute-force simulations.
The authors of Ref. [45] provided numerical evidence for the alleviation of barren plateaus by probing the gradient of the Kullback–Leibler (KL) divergence, for systems of up to twenty qubits. For the numerical examples considered, they found that random initializations led to the gradient variance decaying exponentially with system size (a key signature of the barren-plateau problem), whereas with MPS-pretrained states, the gradient variance did not decay significantly with either system size or the number of gate layers (see Figure 3 in [45] and the related discussion in the main text of [45]).
Furthermore, they extended these findings up to 100 qubits by estimating the gradient magnitude with an MPS-based quantum circuit simulator (for practical feasibility) again with respect to the KL divergence loss. They again found that random initialization led to the gradient vanishing with system size, whereas the gradient magnitude in the pretrained cases exhibited more stable behavior (see Figure 4 in [45] and the related discussion in the main text of [45]).
Overall, these prior findings suggest that the classical computational resources required to bypass the barren-plateau problem do not grow exponentially with system size. Nevertheless, we emphasize that strict scalability is not a prerequisite for achieving a practical advantage. Even if the classical resources needed to avoid barren plateaus become prohibitive at very large system sizes, this approach can still be useful. In this case, the pretraining algorithm may provide a good approximate initialization, which can then be refined on quantum hardware to yield improved performance.
@@ IV.1 Purification and convex combination without pretraining
Prior to training, we initialized the parameters $\eta$, $\nu$, and $\theta$. We set the parameters $\eta$ and $\nu$ in (13) to be equal to zero and one, respectively. We selected the components of the initial parameter vector $\theta$ for the state $\omega(\theta)$ uniformly at random from $0,2\pi]$. The optimal state $\omega$ is almost always a full-rank mixed state. Thus, we used either a three-layer purification ansatz or a convex-combination ansatz with a two-layer quantum circuit Born machine and a two-layer unitary. The particular structure of the purification ansatz and the quantum circuit Born machine we use can be found in a Figure 2 of our companion paper [16].
@@
The variation displayed in the graphs comes from three sources of randomness: the randomized initializations of all parametrized quantum circuits, randomness in the SPSA algorithm for gradient estimation, and shot noise over $10^{12}$ shots for every measurement of a quantum circuit. In our simulations, we were able to achieve an error of order $10^{-2}$ after $20,000$ iterations of training. These results give evidence that using VQE and dual-VQE allows us to bound the true value from both above and below, respectively, as desired.
@@ IV.2 MPS pretraining with purification ansatz
Figure 3: Unpenalized objective function value and penalty value for the Ising model problem using dual-VQE and MPS pretraining. After 2,000 iterations of pretraining followed by 25,000 iterations of quantum training, the true solution is reached when $\chi_{\max}=8$, with a relative error of 0.5%.
@@
The results of the different experiments where we used MPS pretraining are given in Figure 3. The plots shown are for $n=3$ qubits. To demonstrate the effectiveness of pretraining, we used three different max bond dimension values $\chi_{\operatorname{max}}=\{2,4,8\}$ and compared it with the no-pretraining case. In each of the above cases, we conducted five runs and report the mean and standard deviation.
@@
Following the pretraining, we translated the MPS into a quantum circuit. We used the AD decomposition to decompose the MPS into three layers of two-qubit unitaries, labeled as Disentangler in Figure 4. We plot the fidelity of translation, using the AD and OD algorithms (defined in Appendix B), in Figure 4.
The circuit we obtained then serves as the starting point for one of the OD algorithm runs. For this run, labeled as Warm Start, we used 2000 iterations and a constant learning rate of $\beta=0.2$ (see (31)). We also ran the OD algorithm on a randomly initialized three layer circuit and trained for 2000 iterations, labeled as Random Start. Finally, we picked the best translated circuit. For $\chi_{\max}=8$, the translation fidelity was around $99.5\%$.
Lastly, we added a single layer of parametrized gates to the translated circuits, as described in Section III.2.3, and all the parameters are trained using a combination of a finite difference method and the SPSA algorithm. For the parameters $\eta$ and $\nu$, we used a finite difference method with a perturbation of $10^{-6}$. For the circuit parameters, we used the SPSA algorithm since it provides an unbiased estimate of the gradient with a runtime complexity that is independent of the number of parameters. For $\eta$ and $\nu$, we initialized the learning rate, denoted by $\beta_{\eta,\nu}$, to be $10^{-3}$ and for the circuit parameters, we initialized the learning rate, denoted by $\beta_{\omega}$, to be $3\times 10^{-3}$. Every 200 iterations, we modified the learning rates and the penalty parameter $c$. For the learning rates, we calculated the slope of the last 200 iterations. If the slope was negative, we set
@@
$\beta_{\eta,\nu}$ $=\max(0.9\beta_{\eta,\nu},3\beta_{\min}).$ (35)
@@
$\beta_{\eta,\nu}$ $=\min(1.02\beta_{\eta,\nu},3\beta_{max}).$ (37)
@@ Appendix A Matrix product states
Matrix product states have a linear architecture, with each tensor being connected to their nearest neighbors. For example, the tensor $A^{\psi}$ can be factorized into a linear chain of smaller tensors, as seen in Figure 6.
@@
where we use the Einstein repeated index summation convention. The indices labeled by $b_{k}$ are called bonds, and their dimension is called the bond dimension of the MPS. Note that, in general, the bonds can be of different dimension, in which case, the largest bond is quoted as the bond dimension of the MPS and is denoted by $\chi$. In addition to the bond dimension, the size of the output indices is called the physical dimension. In this work, since we are working with qubits, we fix the physical dimension to be $2$. Thus, the number of parameters of the entire MPS is given by $2n\chi^{2}$, as opposed to $2^{n}$ for the full state vector.
@@
The advantages of using tensor networks do not end here. Matrix product operators generalize unitary operations and can be implemented using tensor contractions in a runtime polynomial in the bond dimension. For states with entanglement greater than the max bond dimension can exactly reproduce, the singular value decomposition may be used to find a low bond dimension MPS approximation [47].
Just as mixed states in a Hilbert space can be represented by a pure state in a larger Hilbert space, so too can larger rank MPSs be used to represent mixed states on a smaller rank. For example, consider a mixed state $\omega_{S}$ that has a purification $|\psi\rangle_{RS}$. If $|\psi\rangle_{RS}$ has an MPS representation, then $\omega_{S}$ can be realized as a matrix product operator (MPO) using this MPS, as depicted in Figure 1a).